OpenStudy (anonymous):

can someone check my work? -3x^2 -4x-4=0 (quadratic formula) x=-(-4) sqr -4 - 4(-3)(-4)/2(-3) x=-(-4) sqr 16-48/-6 x=4 sqr 32/-6 x=4 sqr4sqr root of 2/-6 x= 2 sqr 2sqr root of 2/3

OpenStudy (earthcitizen):

first, the equation turns out to be a complex root

OpenStudy (earthcitizen):

\[x = 2/3 (-1-i \sqrt(2))\]

OpenStudy (earthcitizen):

\[x = 2/3 (-1+i \sqrt(2))\]

OpenStudy (anonymous):

ok so the 2i\[\sqrt{2}\] is going to be positive?

OpenStudy (earthcitizen):

yh, that's a complex number with a complex solution

OpenStudy (mertsj):